# Measuring Computational Universality of Fully Homomorphic Encryption

### v1

@@ 4. FHE Libraries Supporting Universality
Table 2 summarizes FHE libraries alongside the encryption schemes they implement and the types of supported operations. In general, HELib, Lattigo, and SEAL offer word-wise FHEs by supporting BFV, BGV and CKKS. In contrast, TFHE and Zama’s TFHE-rs are designed primarily for bit-wise FHEs, offering Boolean gate-level primitives that can be composed into more complex functions. Notably, OpenFHE and Zama extend this universality further by facilitating Scheme Switching or hybrid approaches that bridge bit-wise and word-wise computations. Additionally, methods Polynomial Interpolation (Narumanchi et al., 2017) and Polynomial Interpolation with Special Encoding (Iliashenko and Zucca, 2021) are open-sourced based on HELib, while Encoding Switching (Zhang et al., 2023) is open-sourced using HELib. Functional bootstrapping techniques (Alexandru et al., 2024; Lee et al., 2024) are implemented on Lattigo and OpenFHE but have not yet been open-sourced.
HELib. The Homomorphic Encryption Library (HELib) was introduced in 2013 by IBM and supports the BGV scheme (with bootstrapping), as well as CKKS. Polynomial Interpolation (Narumanchi et al., 2017; Iliashenko and Zucca, 2021) with BGV and Encoding Switching (Zhang et al., 2023) are implemented on HELib and have been open-sourced. HELib is written in C++17 and uses the NTL mathematical library.
Lattigo. The lattice-based multiparty homomorphic encryption library in Go was first developed by the Laboratory for Data Security (LDS) at EPFL and is currently maintained by Tune Insight. It supports BFV, BGV, and CKKS. Lattigo enables scheme switching to compute non-linear functions. Functional FV bootstrapping (Lee et al., 2024) is built on the lattigo, but their implementation is not open-sourced yet.
SEAL. The Simple Encrypted Arithmetic Library (SEAL) was developed by Microsoft Research and was first released in 2015 (mir, 2022b). SEAL supports leveled BFV, BGV, and CKKS.
TFHE. The Fast Fully Homomorphic Encryption Library over the Torus (TFHE) was released in 2016 by Chillotti et al. (Chillotti et al., 2020) and proposes the CGGI cryptosystem. The library exposes homomorphic Boolean gates such as AND and XOR but does not build complex functional units (e.g., adders, multipliers, and comparators) and leaves that to the developer.
OpenFHE. OpenFHE (Al Badawi et al., 2022) is developed by Duality, NJIT, MIT, and other organizations. It supports a wide range of FHE methods, including BGV, BFV, and CKKS with approximate bootstrapping, as well as DM/FHEW, CGGI/TFHE, and LMKCDEY for evaluating Boolean circuits. OpenFHE enables scheme switching between CKKS and between CKKS and FHEW/TFHE to evaluate non-smooth functions, e.g., comparison, using FHEW/TFHE functional bootstrapping. Recently, Alexandru et al. (Alexandru et al., 2024) leveraged OpenFHE to build general functional bootstrapping to enable the non-linear function for any RLWE ciphertexts, but it is not open-sourced yet.
Zama. Zama (Zama, 2022a) is an open-source cryptography company developing state-of-the-art FHE solutions for blockchain and AI. Its products include TFHE-rs (Zama, 2022e) for Boolean and small integer arithmetic, Concrete (Zama, 2022c) for compiling Python to FHE with LLVM, Concrete ML (Zama, 2022b) for encrypted machine learning, and fhEVM for confidential smart contracts in Solidity.
@@
5. Universality Required Privacy-Sensitive Workloads
@@ 5.1. Design Principle
To comprehensively evaluate and compare different universal FHE methods, we designed three workloads requiring universality, each representing basic computational units in privacy-sensitive applications, covering diverse scenarios:
@@ 5.2. Experimental Setup
To evaluate the performance of universal FHE methods, we conducted experiments on the above workloads using input vectors of 100 integers. Comparisons were performed on pairs of these vectors to yield an output vector of 100 Boolean values, while element-wise multiplications on two vectors produced a vector of 100 integers.
@@
System Setup. The experiments are conducted on a server equipped with an AMD Ryzen Threadripper PRO 3955WX (2.2 GHz) and 125 GB of RAM. All tests are run in single-thread mode for a fair comparison. For Encoding Switching, we use the official open-sourced implementation from HEBridge (Zhang et al., 2023) built on HELib. Scheme Switching is adopted from OpenPEGASUS (Lu et al., 2021) built on Microsoft SEAL, and TFHE-rs (Zama, 2022e) are used for the bit-wise TFHE method. All encryption parameters are configured to maintain a security level above 128 bits, following the "BKZ-beta" classical cost model from the LWE estimator (Albrecht et al., 2015) unless stated otherwise.
@@
Parameter Setup. For both BGV/BFV-based methods, i.e., Encoding Switching and Scheme Switching, we set the secret distribution to be a Hamming weight distribution over the set of ternary polynomials with coefficients in $\{0,1,-1\}$, such that each secret has exactly $h=64$ nonzero entries.
@@
For Encoding Switching, which is built on HELib, the plaintext modulus in the FV space is given by $p^{r}$. We choose the parameters $p$ and $r$ based on the input bit-width. Specifically, we use the pairs $\{(4,4),(5,4),(7,5),(17,4)\}$ as $(p,r)$ for input bit-widths of $6$, $8$, $12$, and $16$, respectively. The multiplicative depth is determined by the ciphertext capacity, defined as $\log_{2}\frac{q}{\eta}$, where $q$ is the ciphertext modulus and $\eta$ is the current noise bound. Accordingly, we set $\log_{2}q$ to $\{256,320,488,648\}$ for $6$, $8$, $12$, and $16$ bits of input. After choosing $p$, $r$, and $q$, we select an appropriate degree for the polynomial modulus to ensure that the security level satisfies $\lambda>120$. In this context, the cyclotomic order of the polynomial ring $m$ is chosen so that the ring degree $n$ and the order of the base prime, $d=\texttt{Ord}(p)$, meet the following values: $\{(m,n)=(13201,12852),(16151,15600),(25301,25300),(31621,31212)\}$, corresponding to $6$, $8$, $12$, and $16$ bits of input, respectively.
@@
For Scheme Switching, which is implemented using Microsoft SEAL, we set the parameters in a way that mimics the above setup to ensure a fair comparison. Specifically, we explicitly configure the plaintext modulus $p^{r}$ using the identical $(p,r)$ pairs, and we select the ciphertext modulus $q$ with the same $\log_{2}q$. Then, we choose a polynomial modulus degree $n$ (which in SEAL must be a power of two) such that the overall security level satisfies $\lambda>120$. Although Microsoft SEAL does not allow explicit setting of the cyclotomic order $m$ and the order of $p$ (denoted as $d$), these are implicitly determined by our choices of $n$ and $p$ (with $m$ being typically interpreted as $2n$ for the cyclotomic polynomial $x^{n}+1$, and $d=\texttt{Ord}(p)$ accordingly).
@@
For the bit-wise TFHE method, TFHE-rs library provides objects FheUint6, FheUint8, FheUint12 and FheUint16 to represent differnt input bit precision.
@@ 5.3. Experimental Results
The experimental results are presented in Figure 2. The TFHE (Chillotti et al., 2020), which evaluates both linear and non-linear operations in a bit-wise manner, handles non-linear operations efficiently but performs sub-optimally for linear operations, especially for larger input bit lengths, due to increased circuit depth. More importantly, TFHE lacks SIMD support, a key advantage of word-wise FHE methods that can process multiple ciphertexts simultaneously.
Scheme Switching (Boura et al., 2020), on the other hand, evaluates linear operations using word-wise FHE and non-linear operations using bit-wise FHE. However, the switching process’s complexity grows exponentially with the input bit-width. For instance, on workload-1, evaluating 6-bit inputs takes 9.87 seconds, while 8-bit inputs require 32.1 seconds. This exponential growth makes scheme switching impractical for larger inputs, as similar prohibitive runtimes were estimated for 12-bit and 16-bit inputs.
In contrast, Encoding Switching (Zhang et al., 2023) avoids the overhead of expensive scheme switching and leverages SIMD capabilities by performing both linear and non-linear operations in word-wise FHE. For 8-bit inputs, evaluating workload-1 takes only 15.5 seconds. More importantly, Encoding Switching scales well to larger inputs, taking 23.0 seconds for 12-bit inputs and 46.9 seconds for 16-bit inputs, yielding 14.1$\times$ and 25.4$\times$ improvements over the bit-wise TFHE and Scheme Switching, respectively.
Figure 2. Compare all comparisons on three workloads.
6. Universality Required Privacy-Sensitive Applications
@@ 6.1. Design Principle
In this section, we evaluate universal FHE methods on five representative privacy-sensitive applications that require universal capabilities: private graph algorithms, private decision tree inference, private database aggregation, private neural network inference, and private sorting. These applications cover a diverse range of real-world use cases across domains such as healthcare, finance, and other industrial sectors. Each application incorporates both linear and non-linear operations and is built on the workloads analyzed in Section 5.
@@ 6.2. Experimental Setup
Generally, we evaluate various applications using Scheme Switching, Encoding Switching, and bit-wise TFHE, with experimental setup details identical to those described in Section 5.2. For private graph evaluation, a graph with $n$ nodes is represented as an $n\times n$ adjacency matrix. In the case of bit-wise TFHE, each element of the matrix is individually encrypted into multiple ciphertexts (for multiple bits), whereas for SIMD-enabled methods, each row is batched into a single ciphertext. For private decision tree evaluation, each node is encrypted as multiple ciphertexts for bit-wise FHE, while for SIMD-enabled methods, an entire layer is batched into one ciphertext. In private sorting, each element is encrypted into a single ciphertext as described in (Iliashenko and Zucca, 2021). For private database aggregation, every element is encrypted into multiple ciphertexts when using bit-wise TFHE, whereas for SIMD-enabled methods, each column is batched into a single ciphertext. Finally, for private neural network inference, each parameter is encrypted into multiple ciphertexts with bit-wise TFHE, while for SIMD-enabled methods, each filter or weight matrix is batched into a single ciphertext.
@@ 6.3. Private Graph
Private graph algorithms (Xie and Xing, 2014; Dockendorf et al., 2022; Meng et al., 2015; Li, 2021) enable secure computations on graph-structured data while preserving individual privacy (Xie and Xing, 2014; Mittal et al., 2012; Akcora et al., 2012). In this threat model, an untrusted server performs computations without learning any details about the input graph. A common operation in graph algorithms is selecting the shorter path from paths $P_{1}$ and $P_{2}$, which can be implemented as: $(P_{1}>P_{2})\times P_{2}+(1-(P_{1}>P_{2}))\times P_{1}$; this computation exemplifies workload-2, where a linear operation follows a non-linear operation. In this section, we demonstrate the performance of universal FHE methods on the private Floyd-Warshall algorithm (Floyd, 1962).
@@
Bit-wise FHE Private Floyd-Warshall. To implement a private Floyd-Warshall algorithm, a naive method (Dockendorf et al., 2022) encrypts each element of the adjacency matrix $G$ to ciphertext. The server then conducts the same procedure on the encrypted $G$ as delineated in Algorithm 1. For if-else operations based on comparison results, the server uses HE multiplication to implement the conditional logic. Specifically, lines 10-11 of Algorithm 1 can be computed in HE as:
@@
Figure 3 illustrates this SIMD optimization for a graph with $4$ nodes, focusing on updating distances from node $1$ to all others through node $4$. The process is: ❶ Generate ciphertext $\bm{D_{1\rightarrow 4}}$ by multiplying $[D_{1\rightarrow 1},D_{1\rightarrow 2},D_{1\rightarrow 3},D_{1\rightarrow 4}]$ with plaintext $[0,0,0,1]$, then replicating $D_{1\rightarrow 4}$ across all slots by rotations and addition. ❷ Compute $\bm{D_{1\rightarrow 4\rightarrow all}}$ by adding $\bm{D_{1\rightarrow 4}}$ to $\bm{D_{4\rightarrow all}}$, representing new potential distances via node $4$ Compare $\bm{D_{1\rightarrow 4\rightarrow all}}$ and $\bm{D_{1\rightarrow all}}$, yielding mask $\bm{M}$ ($1$ where new distance is shorter, 00 otherwise). ❹ Update distances by:
(12) $\bm{D_{1\rightarrow all}}=\bm{M}\cdot\bm{D_{1\rightarrow 4\rightarrow all}}+(1-\bm{M})\cdot\bm{D_{1\rightarrow all}},$
only paths via node $4$ is shorter than original one will be updated.
@@
9:        $D_{i\rightarrow k}=\text{SIMD}(D[i,k])$ Broadcast $D[i,k]$ to all slots
@@
12:        $M=D_{\text{new}}<D[i,:]$ Mask for minimum
@@
Experimental Analysis. To assess the performance of different universal FHE methods for the Private Floyd-Warshall algorithm, we conducted experiments across various input bit widths and graph sizes. Figure 4 (a) compares the performance for a graph with 32 nodes under different input bit lengths, where Encoding Switching exhibits the best performance and Scheme Switching the worst, particularly for larger bit widths. Figure 4 (b) shows results for graphs of varying sizes using an 8-bit input. As noted earlier, the runtime of bit-wise TFHE grows approximately as $O(|V|^{3})$ due to the lack of SIMD support, while the complexities for both Scheme Switching and Encoding Switching are reduced to roughly $O(|V|^{2})$. For 8-bit inputs, TFHE takes 13.3 seconds, and Encoding Switching takes 17.3 seconds on a 16-node graph; however, for a 128-node graph, TFHE’s runtime escalates dramatically to 4,933 seconds, whereas Encoding Switching completes in 632 seconds.

### v2

@@ 5. FHE Libraries Supporting General Computation
Table 3 summarizes FHE libraries alongside the encryption schemes they implement and the types of supported operations. In general, HELib, Lattigo, and SEAL offer word-wise FHEs by supporting BFV, BGV and CKKS. In contrast, TFHE and Zama’s TFHE‑rs are designed primarily for bit-wise FHEs, offering Boolean gate-level primitives that can be composed into more complex functions. Notably, OpenFHE and Zama extend this generality further by facilitating Scheme Switching or hybrid approaches that bridge bit-wise and word-wise computations. Additionally, methods Polynomial Interpolation (narumanchi2017performance) and Polynomial Interpolation with Special Encoding (iliashenko2021faster) are open-sourced based on HELib, while Encoding Switching (zhang2023hebridge) is open-sourced using HELib. Functional bootstrapping techniques (alexandru2024general; lee2024functional) are implemented on Lattigo and OpenFHE but have not yet been open-sourced.
@@
HELib. The Homomorphic Encryption Library (HELib) was introduced in 2013 by IBM and supports the BGV scheme (with bootstrapping), as well as CKKS. Polynomial Interpolation (narumanchi2017performance; iliashenko2021faster) with BGV and Encoding Switching (zhang2023hebridge) are implemented on HELib and have been open-sourced. HELib is written in C++17 and uses the NTL mathematical library.
Lattigo. The lattice-based multiparty HE library in Go was first developed by the Laboratory for Data Security (LDS) at EPFL and is currently maintained by Tune Insight. It supports BFV, BGV, and CKKS. Lattigo enables scheme switching to compute non-linear functions. Functional FV bootstrapping (lee2024functional) is built on the lattigo, but their implementation is not open-sourced yet.
SEAL. The Simple Encrypted Arithmetic Library (SEAL) was developed by Microsoft Research and was first released in 2015 (mircroSEAL). SEAL supports leveled BFV, BGV, and CKKS.
TFHE. The Fast Fully Homomorphic Encryption Library over the Torus (TFHE) was released in 2016 by Chillotti et al. (chillotti2020tfhe) and proposes the CGGI cryptosystem. The library exposes homomorphic Boolean gates such as AND and XOR, but does not build complex functional units (e.g., adders, multipliers, and comparators) and leaves that to the developer.
OpenFHE. OpenFHE (al2022openfhe) is developed by Duality, NJIT, MIT, and other organizations. It supports a wide range of FHE methods, including BGV, BFV, and CKKS with approximate bootstrapping, as well as DM/FHEW, CGGI/TFHE, and LMKCDEY for evaluating Boolean circuits. OpenFHE enables scheme switching between CKKS and between CKKS and FHEW/TFHE to evaluate non-smooth functions, e.g., comparison, using FHEW/TFHE functional bootstrapping. Recently, Alexandru et al. (alexandru2024general) leveraged OpenFHE to build general functional bootstrapping to enable the non-linear function for any RLWE ciphertexts, but it is not open-sourced yet.
Zama. Zama (Zama) is an open-source cryptography company developing state-of-the-art FHE solutions for blockchain and AI. Its products include TFHE-rs (TFHE-rs) for Boolean and small integer arithmetic, Concrete (Concrete) for compiling Python to FHE with LLVM, Concrete ML (ConcreteML) for encrypted machine learning, and fhEVM for confidential smart contracts in Solidity.
6. General Computation Required Workloads
@@ 6.1. Design Principle
To comprehensively evaluate and compare different general FHE methods, we designed three workloads requiring generality, each representing basic computational units in privacy-sensitive AI applications, covering diverse scenarios:
@@ 6.2. Experimental Setup
To evaluate the performance of general FHE methods, we conducted experiments on the above workloads and measured both the total execution time and the amortized time for SIMD-enabled word-wise Scheme switching and Encoding switching.
@@
System Setup. The experiments are conducted on a server equipped with an AMD Ryzen Threadripper PRO 3955WX (2.2 GHz) and 125 GB of RAM. All tests are run in single-thread mode for a fair comparison. For Encoding Switching, we use the official open-sourced implementation from HEBridge (zhang2023hebridge) built on HELib. Scheme Switching is adopted from OpenPEGASUS (lu2021pegasus) built on Microsoft SEAL, and TFHE-rs (TFHE-rs) are used for the bit-wise TFHE method. All encryption parameters are configured to maintain a security level above 128 bits (lauter2022protecting; bossuat2024security), following the ”BKZ-beta” classical cost model from the LWE estimator (albrecht2015concrete) unless stated otherwise.
@@
Parameter Setup. Parameter choices ensure operations remain within the multiplicative depth budget, avoiding word-wise bootstrapping. For both BGV/BFV‐based methods, i.e., Encoding Switching and Scheme Switching, we set the secret distribution to be a Hamming weight distribution over the set of ternary polynomials with coefficients in $\{0,1,-1\}$, such that each secret has exactly $h=64$ nonzero entries.
@@
For Encoding Switching, which is built on HELib, the plaintext modulus in the FV space is given by $p^{r}$. We choose the parameters $p$ and $r$ based on the input bit-width. Specifically, we use the pairs $\{(4,4),(5,4),(7,5),(17,4)\}$ as $(p,r)$ for input bit-widths of $6$, 88, $12$, and $16$, respectively. The multiplicative depth is determined by the ciphertext capacity, defined as $\log_{2}\frac{q}{\eta}$, where $q$ is the ciphertext modulus and $\eta$ is the current noise bound. Accordingly, we set $\log_{2}q$ to $\{256,320,488,648\}$ for $6$, 88, $12$, and $16$ bits of input. After choosing $p$, $r$, and $q$, we select an appropriate degree for the polynomial modulus to ensure that the security level satisfies $\lambda>128$. In this context, the cyclotomic order of the polynomial ring $m$ is chosen so that the ring degree $n$ and the order of the base prime, $d=\texttt{Ord}(p)$, meet the following values: $\{(m,n)=(13201,12852),(16151,15600),(25301,25300),(31621,31212)\}$, corresponding to $6$, 88, $12$, and $16$ bits of input, respectively.
@@
For Scheme Switching, which is implemented using Microsoft SEAL, we set the parameters in a way that mimics the above setup to ensure a fair comparison. Specifically, we explicitly configure the plaintext modulus $p^{r}$ using the identical $(p,r)$ pairs, and we select the ciphertext modulus $q$ with the same $\log_{2}q$. Then, we choose a polynomial modulus degree $n$ (which in SEAL must be a power of two) such that the overall security level satisfies $\lambda>128$. Although Microsoft SEAL does not allow explicit setting of the cyclotomic order $m$ and the order of $p$ (denoted as $d$), these are implicitly determined by our choices of $n$ and $p$ (with $m$ being typically interpreted as $2n$ for the cyclotomic polynomial $x^{n}+1$, and $d=\texttt{Ord}(p)$ accordingly).
@@
For the bit-wise TFHE method, TFHE-rs provides FheUint6, FheUint8, FheUint12, FheUint16, FheUint24, and FheUint32 to represent differnt input bit precision.
@@ 6.3. Experimental Results
Figure 2(a) reports the latency of a single run, whereas Figure 2 (b) illustrates the amortized time across slots for FHE methods supporting SIMD parallelism (Scheme Switching and Encoding Switching).
TFHE (chillotti2020tfhe) evaluates all operations bit-wise, handling non-linear functions efficiently but performing poorly on linear ones at larger bit lengths. It lacks SIMD support, unlike word-wise schemes that process multiple ciphertexts simultaneously. Notably, TFHE performs better on workload-2 than workload-1 because, in workload-2, one multiplication input is the comparison result (all 0s or all 1s), reducing circuit depth compared to workload-1, where multiplication precedes comparison
Scheme Switching (boura2020chimera), on the other hand, evaluates linear operations using word-wise FHE and non-linear operations using bit-wise FHE. However, the switching process’s complexity grows exponentially with the input bit-width. For instance, on workload-1, 6-bit inputs take 9.87 seconds and 8-bit inputs 32.1 seconds, with runtimes for 12- and 16-bit inputs becoming prohibitive, making the approach impractical for larger inputs.
[…]
When comparing the amortized time in Figure 2 (b) with the single-run results, TFHE overtakes scheme switching as the slowest method, owing to its lack of SIMD support.
Table 4 reports storage and communication metrics for the three general FHE methods on Workload-1 ($1\times 100$ vector of 8-bit integers). TFHE achieves the smallest footprint (4.8 MB ciphertext, 9.6 MB communication, 23.3 MB memory), making it ideal for bandwidth-constrained scenarios, while Encoding Switching exhibits 10× larger overhead (47.6 MB, 95.2 MB, 94.3 MB) but enables faster computation in resource-rich environments. These metrics introduce deployment constraints into method selection: applications with limited bandwidth or storage should prioritize TFHE despite higher computational costs, whereas datacenter deployments can accept Encoding Switching’s larger footprint for performance gains.
Table 4. Storage and communication overhead (Workload-1).
Method Ciphertext Peak Memory Communication
TFHE 4.8 MB 23.3 MB 9.6 MB
Scheme Switching 14.2 MB 58.6 MB 28.4 MB
Encoding Switching 47.6 MB 94.3 MB 95.2 MB
Communication between server and client.
7. General Computation Required Applications
@@ 7.1. Design Principle
In this section, we evaluate general FHE methods on five representative privacy-sensitive AI applications that require general computation: private graph algorithms, private decision tree inference, private database aggregation, private neural network inference, and private sorting. These applications cover a diverse range of real-world use cases across domains such as healthcare, finance, and other industrial sectors. Each application incorporates both linear and non-linear operations and is built on the workloads analyzed in Section 6.
@@ 7.2. Experimental Setup
Generally, we evaluate various applications using Scheme Switching, Encoding Switching, and bit-wise TFHE, with experimental setup details identical to those described in Section 6.2. For private graph evaluation, a graph with $n$ nodes is represented as an $n\times n$ adjacency matrix. In the case of bit-wise TFHE, each element of the matrix is individually encrypted into multiple ciphertexts (for multiple bits), whereas for SIMD-enabled methods, each row is batched into a single ciphertext. For private decision tree evaluation, each node is encrypted as multiple ciphertexts for bit-wise FHE, while for SIMD-enabled methods, an entire layer is batched into one ciphertext. In private sorting, each element is encrypted into a single ciphertext as described in (iliashenko2021faster). For private database aggregation, every element is encrypted into multiple ciphertexts when using bit-wise TFHE, whereas for SIMD-enabled methods, each column is batched into a single ciphertext. Finally, for private neural network inference, each parameter is encrypted into multiple ciphertexts with bit-wise TFHE, while for SIMD-enabled methods, each filter or weight matrix is batched into a single ciphertext.
@@ 7.3. Private Graph
Private graph algorithms (xie2014cryptgraph; dockendorf2022graph; meng2015grecs; li2021checking) enable secure computations on graph-structured data while preserving individual privacy (xie2014cryptgraph; mittal2012preserving; akcora2012privacy). In this threat model, an untrusted server performs computations without learning any details about the input graph. A common operation in graph algorithms is selecting the shorter path from paths $P_{1}$ and $P_{2}$, which can be implemented as: $(P_{1}>P_{2})\times P_{2}+(1-(P_{1}>P_{2}))\times P_{1}$; this computation exemplifies workload-2, where a linear operation follows a non-linear operation. In this section, we demonstrate the performance of general FHE methods on the private Floyd-Warshall algorithm (floyd1962algorithm).
@@
Bit-wise FHE Private Floyd-Warshall. To implement a private Floyd-Warshall algorithm, a naive method (dockendorf2022graph) encrypts each element of the adjacency matrix $G$ to ciphertext. The server then conducts the same procedure on the encrypted $G$ as delineated in Algorithm 1. For if-else operations based on comparison results, the server uses HE multiplication to implement the conditional logic. Specifically, lines 10-11 of Algorithm 1 can be computed in HE as:
@@
Figure 3 illustrates this SIMD optimization for a graph with $4$ nodes, focusing on updating distances from node $1$ to all others through node $4$. The process is: ❶ Generate ciphertext $\boldsymbol{D_{1\rightarrow 4}}$ by multiplying $D_{1\rightarrow 1},D_{1\rightarrow 2},D_{1\rightarrow 3},D_{1\rightarrow 4}]$ with plaintext $[0,0,0,1]$, then replicating $D_{1\rightarrow 4}$ across all slots by rotations and addition. ❷ Compute $\boldsymbol{D_{1\rightarrow 4\rightarrow all}}$ by adding $\boldsymbol{D_{1\rightarrow 4}}$ to $\boldsymbol{D_{4\rightarrow all}}$, representing new potential distances via node $4$. ❸ Compare $\boldsymbol{D_{1\rightarrow 4\rightarrow all}}$ and $\boldsymbol{D_{1\rightarrow all}}$, yielding mask $\boldsymbol{M}$ (11 where new distance is shorter, 0 otherwise). ❹ Update distances by:
(2) $\boldsymbol{D_{1\rightarrow all}}=\boldsymbol{M}\cdot\boldsymbol{D_{1\rightarrow 4\rightarrow all}}+(1-\boldsymbol{M})\cdot\boldsymbol{D_{1\rightarrow all}},$
Only paths via node $4$ that are shorter than the original one will be updated.
@@
9:    $D_{i\rightarrow k}=\text{SIMD}(D[i,k])$ # Broadcast $D[i,k]$ to all slots
@@
12:    $M=D_{\text{new}}<D[i,:]$ # Mask for minimum
@@
Experimental Analysis. To assess the performance of different general FHE methods for the Private Floyd-Warshall algorithm, we conducted experiments across various input bit widths and graph sizes. Figure 4 (a) compares the performance for a graph with 32 nodes under different input bit lengths, where Encoding Switching exhibits the best performance and Scheme Switching the worst, particularly for larger bit widths. Figure 4 (b) shows results for graphs of varying sizes using an 8-bit input. As noted earlier, the runtime of bit-wise TFHE grows approximately as $O(|V|^{3})$ due to the lack of SIMD support, while the complexities for both Scheme Switching and Encoding Switching are reduced to roughly $O(|V|^{2})$. For 8-bit inputs, TFHE takes 14.69 seconds, and Encoding Switching takes 17.28 seconds on a 16-node graph; however, for a 128-node graph, TFHE’s runtime escalates dramatically to 5,432 seconds, whereas Encoding Switching completes in 632 seconds.